OpenStudy (anonymous):

Find the least common denominator for these two rational expressions. -4/x^2 and 6/9x

OpenStudy (anonymous):

help meee

OpenStudy (anonymous):

sorry it should by 3x^2

OpenStudy (anonymous):

no the other answer was right

OpenStudy (anonymous):

it is right but could be simplified

OpenStudy (anonymous):

6/9x=(2*3)/(3*3*x)=2/3x

OpenStudy (anonymous):

thank you can u help me with the other question

OpenStudy (anonymous):

ok